In this problem, y = 1/(x2 + c) is a one-parameter family of solutions of the first-order
DE y' + 2xy2 = 0.Find a solution of the first-order IVP consisting of this differential equation and the given initial condition.

y(3) = 1/5

y = ___

Give the largest interval I over which the solution is defined.

A (0, infinity)
B. (2, infinity)
C. ( - infinity, infinity)
D. (0, 2)
E. (-2, infinity)
1
Expert's answer
2020-06-17T18:57:48-0400

Given,

"y=\frac{1}{x^2+c}"

is family solution of

"y'+2xy^2=0"

Initial conditions is "y(3)=\frac{1}{5}" ,Thus, from General solution we get,

"\frac{1}{5}=\frac{1}{(3^2+c)}\implies c=-4"

Thus, particular solution will be

"y(x)=\frac{1}{x^2-4}"

Clearly, "y" is not defined at "x=\pm 2" , thus larges interval in which solution is defined is

"(-\infty,\infty)\backslash\{-2, 2\}"

Acceptable given option will be (B).